### v1

# Victoria, BC, Canada

Abstract
Mixed and multi methods research is often used in software engineering, but researchers outside of the social or human sciences often lack experience when using these designs. This paper provides guidelines and advice on how to design mixed and multi method research, and to encourage the intentional, rigourous, and innovative use of mixed methods in software engineering. It also presents key characteristics of core mixed method research designs. Through a number of fictitious but recognizable software engineering research scenarios and personas of prototypical researchers, we showcase how to choose suitable designs and consider the inevitable tradeoffs any design choice leads to. We furnish the paper with recommended best practices and several antipatterns that illustrate what to avoid in mixed and multi method research.
Keywords:
Mixed methods Multi methods Research methods Methodology Guidelines
1 Introduction
Software engineering (SE) researchers use a variety of research methods, often mixing and matching them. Despite the widespread prevalence of mixing methodsbryman_integrating_2006 ; creswell_best_2011 ; poth_innovation_2018 ; ladner_mixed_2019 ; poth_sage_2023 , there is little guidance in SE on how to do so systematically, in a way that is well motivated, carefully designed, robustly executed, and ideally, leading to additional benefits not otherwise achievable from the use of a single research method alone. Mixing methods is particularly pertinent in today’s complex socio-technical SE contexts where multiple methods can bring insights on the nuanced nature of the problem or solution being studied.
Our paper is aimed at researchers unfamiliar with mixed methods. We share key concepts and provides guidance in how to mix and match methods, presenting different characteristics of mixed methods and some common designs. Inspired by our previous paper on selecting empirical methods for software engineering research easterbrook_selecting_2008 , this paper provides a number of illustrative fictitious research scenarios that capture different research aims and call for very different research design choices. These scenarios help to guide design choice but also emphasize the tradeoffs faced in making mixed method study design decisions. As well as showcasing design choices for a variety of scenarios, the paper also calls out best practices and a number of antipatterns to be avoided.
The timing of this paper is important, as budding SE researchers today face a complex world where software is used to address complex problems with wide-scale societal impact. There is also a strong drive for innovation in the face of rapidly changing technologies and practices, all occurring in socio-technical contexts storey_who_2020 ; hoda_socio-technical_2021 . The increasing interest in studying the adoption of generative AI technologies in everyday SE practice is but one latest example of the constantly changing SE research landscape bano2023LLM . Choosing which mixed methods to use and knowing how to use them effectively is challenging, and is often overlooked in SE research and education curricula and thus, experience is lacking in how to use these methodological designs.
In this paper, we include practical illustrative examples in the form of four fictional SE research personas and scenarios, representing common mixed method contexts, designs, and outcomes. These four personas and their research scenarios are used to exemplify and motivate the use of mixed and multiple method research designs throughout the paper. The personas we introduce include:
  • Ali, a postdoc studying how the addition of security experts to development teams impacts developer motivations and behaviours;
  • Sam, an early stage Ph.D. student studying the impact of a new code review tool at a small company;
  • Vicki, an experienced industry researcher studying the adoption of generative AI at their company; and
  • Zara, a late stage Ph.D. student evaluating an automated tool they designed.
The rest of this paper is organized as follows. In Section 2, we present a brief background of the most prominent and relevant influences in mixed and multi methods research design over the decades. In Section 3, we present four core principles that underlie mixed method research: methodological rationale, novel insights, procedural rigor and ethical research. In Section 4, we discuss some notable characteristics of mixed method designs, as well as common mixed and multi method research designs. In Section 5, we present the four personas of budding new researchers in software engineering and detail the challenging research decisions they face. We describe the methodological choices they make and how they followed the four core principles mentioned in Section 3. In Section 6, we use the concept of antipatterns to present what not to do in mixed method research before concluding the paper in Section 7. This paper includes an Appendix which contains a glossary of terms we use throughout the paper for ease of reference.
2 A Brief Overview of Mixed Methods Research
Early researchers recognized that many research questions can be addressed using both qualitative (quaL) and quantitative (quaN) methods that are combined in various stages of the investigation process, from data collection and analysis, to synthesis and to the reporting of findings. The need to integrate quaL descriptions and quaN measurements to understand phenomena is not new. Early roots of mixed methods date back as far as 1000 BC in the natural sciences and were used by Babylonian astronomers. Aristotle and Galileo also referred to use of mixed methods for classifying animals and observing planets respectively maxwell_expanding_2016 .
It is only in the mid-20th century, with roots in disciplines such as sociology and anthropology, that MMR gained traction with the formalization of approaches by authors including Creswell and Plano Clark creswell_designing_2018 . According to these authors, “The formative period of mixed methods began in the 1950s and continued up until the 1980s. This period saw the initial interest in using more than one method in a study” but these authors did not explicitly identify these approaches as mixed methods. Furthermore, the different typologies of mixed method designs did not emerge until much later maxwell_expanding_2016 . In the social sciences field, Campbell and Fisk’s paper in 1959campbell_convergent_1959 is recognized as the first adoption of multiple research methods for validation purposes. They also defined this design as multi-operationalism since the authors propose a multitrait, multimethod matrix that intersects quaL-quaN methods with a specific set of traits as a means for enforcing validity of research conclusions. In the mid 1960s, Webb et.al webb_unobtrusive_1966 introduced the concept of triangulation, followed by Denzin in the 1970s denzin_research_1978 who detailed how to triangulate using both within-methods triangulation and between-methods triangulation, emphasizing the latter as a more powerful means to overcome bias and as a way to pragmatically mix methods.
Although the combination of quaL-quaN methods has been used in practice for many years, the first evidence of a more formal use of mixed methods is identifiable only in the late 1980s, discussed by Guest and Fleming guest_mixed_2015 and Johnson et al. johnson_toward_2007 who coined mixed methods research as the third paradigm, coexisting with quaN and quaL studies. The common denominator in these early conceptualizations of mixed methods is that MMR combines quaL-quaN elements. It is only recently that researchers have turned their attention to proposing typologies of mixed methods, i.e., which describe how to structure and design a MMR by combining quaL and quaN elements according to some criteria as stated by Teddlie and Tashakkori teddlie_foundations_2009 .
This rise of formalized mixed method approaches has led to numerous exciting discussions, but often with different design representations and overlapping terms. Indeed, not even the basic definition of what can be qualified as an MMR study sees consensus (see Section 1 in poth_sage_2023 ). Part of this issue originates in the “paradigm war”, where some researchers have polarized preferences for using quaN (e.g., positivism) or quaL (e.g., constructivism) data methods. It is not our goal to review these research paradigms, design typologies or resolve paradigm debates in this paper, but it is essential to note that some research paradigms that underpin MMR, such as pragmatism, emphasize the importance of using the most effective methods to answer research questions. Moreover, the philosophical foundation of a researcher’s choice will inevitably impact how quaL and quaN approaches are integrated. Today, MMR is widely recognized and utilized across diverse fields such as psychology, education, public health, and management, and these approaches offer a flexible and robust framework for addressing multifaceted research inquiries and advancing knowledge in complex socio-technical domains poth_innovation_2018 ; ladner_mixed_2019 .
With the increasing complexity of MMR designs, the problem of discussing and visually presenting complex designs to readers became apparent. Consequently, as noted by Creswell and Plano Clark poth_sage_2023 when they revisited MMR designs twenty years after their important chapter was first published in 2003 tashakkori_handbook_2003 , the most recent studies decreased their focus on the quaN and quaL components and increased attention towards the integration processes across design types. To reflect this change about the central elements of the MMR designs, Figure 1 represents the simplification of the interconnection of the core components of a MMR design. From this perspective, treating the data beyond quaL and quaN analysis and gaining insight through integration and meta-inferences become central.
Figure 1: Interconnection of the core components for MMR: collecting, analyzing and interpreting (1) qualitative and (2) quantitative data to address a research question; (3) integrating the data (merging, connecting, and building) in a research design and set of procedures; and (4) drawing meta-inferences that add new insight beyond the quantitative and qualitative data. Adapted from Creswell and Plano Clark poth_sage_2023 .
To conclude our brief overview of MMR, current trends in MMR can be listed as a call for openness and creativity during the research process, integrating mixed methods in interdisciplinary research, and the use of digital tools for data collection and analysis. Furthermore, researchers will be increasingly challenged to keep pace with a rapid evolution in design terminology and naming practices. Given such dynamism, paying attention to the many sources of contextual influences on the design of MMR becomes paramount poth_sage_2023 . Recognizing the challenges these trends impose on researchers, we provide practical design guidance of MMR for SE researchers in our paper. To start, in the next section, we introduce the definition of MMR we use and then we describe four fundamental principles to be followed when using MMR.
3 Principles to Guide Mixed and Multi Method Research
Figure 2: Principles for Mixed and Multi Method Research Designs
Building on Poth poth_innovation_2018 , we define mixed and multi methods research (MMMR) [^1] Expanding the MMR abbreviation to include an additional ‘M’ for multi. as, A research approach where multiple procedures are used to collect, analyze, and integrate data to address a research problem. For mixed methods, both qualitative and quantitative data are integrated to produce novel insights.
The difference between mixed methods and multi methods lies in what types of data are used and how the data is used. Mixed methods requires the integration of quantitative (quaN) and qualitative (quaL) data. Multi methods can be applied with any combination including quaN-quaN or quaL-quaL.
Poth poth_innovation_2018 also presents four core guiding principles for MMMR that we have adapted slightly to be applicable to software engineering (see Figure 2). Two of the principles, methodological rationale and novel insights, relate to understanding why we should consider using a mixed or multiple methods research design. The other two, procedural rigor and ethical research, relate to understanding how we should go about designing and applying mixed and multi methods research. We delve into these four principles below.
3.1 Methodological Rationale
Principle of Methodological Rationale: MMMR study designs explain how the mixed and multiple research design and data procedures are relevant for addressing a given research problem. An MMMR study description should answer the question why did we select a particular MMMR research design and set of procedures?
There are several conceptual reasons for using mixed or multiple methods [^2] The reasons we list are inspired by or earlier proposed by Ladner ladner_mixed_2019 , Bryman bryman_social_2012 , Creswell creswell_designing_2018 and McGrath mcgrath_methodology_1995 . The rationale for the methodology may be one or more of the following:
  • Complementarity: one procedure helps to deepen, confirm or enhance the insights of another procedure ladner_mixed_2019 ; bryman_social_2012 .
  • Expansion: one procedure is used to ask different questions that are apparent up front or emerge throughout the study (in parallel or sequentially) ladner_mixed_2019 ; bryman_social_2012 .
  • Development: one procedure is used to inform or improve another procedure (e.g., interviews may be used to inform the development of a questionnaire) ladner_mixed_2019 ; bryman_social_2012 .
  • Triangulation: a different type of procedure is used to corroborate earlier insights leveraging multiple quality criteria afforded by use of different procedures mcgrath_methodology_1995 notably generalizability, realism, precision, and control.
  • Credibility: another procedure may be used to increase the integrity and truth of the findings from a former procedure bryman_social_2012 . Multiple or mixed methods can increase credibility by bringing multi-dimensional philosophical and theoretical lenses to the findings mertens_mixed_2013 .
  • Explanation: one procedure can be used to resolve previous contradictory, surprising or inconclusive findings bryman_integrating_2006 .
  • Increased design flexibility: implies researchers may need to consider research design alternatives to offset limitations of a single method when faced with unexpected issues or unsatisfactory data. For example, many researchers that were planning to conduct ethnographic studies during the global pandemic pivoted to using alternative designs, such as a combination of online interviews, self-reported work diary entries, and data mining [^3] https://www.microsoft.com/en-us/research/project/the-new-future-of-work/.
3.2 Novel Insights
Principle of Novel Insights: MMMR studies generate new insights through the integration of the procedures, which in turn addresses the reason(s) for the use of multiple or mixed methods. A well-executed MMMR study answers the question what did we gain from using MMMR?
Succinctly, an MMMR study should deliver on its promise of additional benefits to what could have been afforded by using a single method:
  • Improved problem understanding: a better understanding of the problem space and a fuller picture of the research question before rushing into a solution.
  • Greater depth and breadth: more nuanced and comprehensive coverage of the problem and solution spaces, with room for evaluating the proposed solutions.
  • Explaining unexpected results: where the use of one procedure produced unexpected results, another can be employed in a bid to further explore why those unexpected results were encountered.
  • Complementary storytelling: quaN data provides a numerical strength and statistical measure, and quaL data provides richer insights into the experiences and perspectives of those close to the phenomenon. Together they help narrate a better and fuller story, with multiple ways to present findings in engaging formats [^4] Bryman refers to this as illustration, or the idea of “putting ‘meat on the bones’ of ‘dry’ quaN findings” bryman_social_2012 .
3.3 Procedural Rigor
Principle of Procedural Rigor: Mixed and multi method research studies enhance the reliability and validity of the findings throughout the study. A well-executed MMMR study answers the question how well did we conduct the study?
Rigor in any research study is always a critical concern, and it is often a difficult concept for students to know how to assess. Rigor is particularly challenging for mixed and multi methods. MMMR studies should first adhere to the rigor of the individual studies, but there are also some concerns that apply to the “mixing” aspect. For this principle, we are guided by a mixed methods rigor appraisal tool called MMAT (Mixed Methods Appraisal Tool)  oliveira_mixed_2021 . The following are posed as questions to be considered when self-reflecting on the rigor of a mixed design:
  • Is the use of mixed methods justified? As mentioned above, methodological rationale (e.g., those listed above) should be clearly described.
  • Are the methods effectively integrated to answer the research question? It is important to appraise how well the data from the multiple methods are integrated to answer the research question posed and to appraise at what stage the methods are integrated: during collection or during analysis.
  • Are the findings from the mixed or multiple methods integrated? One must ask what are the similarities and divergencies in the findings (the output of the methods), asking does the integration add to more than the sum of the multiple parts? Any contradictions or divergences should be explained or discussed.
  • Are the different methods used rigorously conducted? MMMR studies should adhere to the rigor of the individual studies used in the mixed or multi study (referring to the appropriate quality criteria for each study).
3.4 Ethical Research
Principle of Ethical Research: MMMR studies manifest the application of responsible research, with respect for people (e.g., participants, concerned parties, beneficiaries) and their welfare (e.g., respect for their confidentiality, privacy, time, emotions, cultural sensitivity, and needs), and that of the environment (e.g., organisations, communities of practice, society, and the planet at large, aiming for sustainability). A good MMMR study answers the question how responsibly did we do it?
Ethical research is a timeless concept. However, the more recent unchecked advancements in AI and unprecedented access to publicly available datasets has meant that SE advancements are progressing faster than ethical standards, guidelines, and regulations can keep up. Both SE industry and research communities are calling for action to take stock of where we are headed collectively in terms of ethical practice in the software industry. Industrial and academic researchers alike need to be held accountable for, be aware of, and proactively manage the ethical aspects of any research work.
An ethical approach to research starts by carefully considering the why and not being driven by the why not. This applies to the use of any procedure, but is even more pertinent to consider with mixed or multi methods because of the additional risks these designs bring in terms of privacy and confidentiality when mixing and integrating different types of data.
  • Considering the why: Just because we may have access to large datasets and the latest in computational advancements, should we rush into designing research that can be performed using them or about them? What are the driving motivations for conducting the research in the first place? What benefits may it bring for the community being studied and for a wider society? Similarly, just because we can design a survey to ask some relevant questions, should we be spamming practitioner mailing lists with requests to fill them out? For example, contacting individuals, contributors, or leaders privately without explicit permission has been deemed as unacceptable behaviour by GitHub [^5] https://github.com/community/community/blob/main/CODE_OF_CONDUCT.md.
  • Privacy and confidentiality: Deep and revealing conversations with research participants lend richness and nuance to the findings. Similarly, access to sensitive data for research purposes, such as team chats, emails, bug reports, and financial reports, is a privilege. What are our plans for protecting participant and data privacy and confidentiality when gathering, storing, and processing these assets, and when presenting findings based on them?
  • Respect and cultural sensitivity: Many studies involve people from multiple locations and cultural backgrounds. Have we considered these cultural sensitivities when designing our protocols? Do we need to customise our protocols to make better sense to different participant groups or be more culturally appropriate in some communities? Respect extends to a participant’s time and privacy, e.g., how will we ensure the privacy of traceable information leading to individual identification in open-source repositories or in end-user reviews of software apps and products? Some types of mining of publicly available data can be considered unintended and an unapproved use, especially where the chain of evidence can lead to identification and loss of privacy.
  • Safety and welfare: A cornerstone of ethical research is to ensure the safety and welfare of the people and environments touched by the research. What are the physical, virtual, or psychological risks we may be exposing our participants and ourselves to? What are the environmental implications of our research? Do we have the requisite training to conduct risk assessments and to ensure the safety of our participants, ourselves, and the environment?
  • Special concerns about MMMR: While many of the ethical concerns above may equally apply to a single method research study, there are some that may be raised specifically due to the use of MMMR. Typically, MMMR studies collect and analyze a variety of data from multiple sources. While this is a strength of MMMR studies, it is also a potential threat. Because of the multiple angles at which an MMMR study approaches a phenomenon, mixing at the data level can place participant privacy and confidentiality under threat. For example, if recruiting survey participants for follow-up interviews from a survey cohort, it may be feasible to trace interviews back to survey responses using demographic data. Similarly, interview responses can be traced back to email, chat, or code commit logs, jeopardising the privacy and confidentiality of certain participants or data sources. Thus, special care needs to be taken when using MMMR.
4 A Landscape of Mixed and Multi Methods Design
Research study design that involves choosing suitable procedures and deciding on their timing is a challenging process easterbrook_selecting_2008 , but research design is particularly challenging in the case of an MMMR study. The research problem or questions being asked and how they emerge over the course of a study guides the research design and the timing of procedures used. Below we discuss the main characteristics of a research design, followed by some common patterns of MMMR designs.
4.1 Core Characteristics of Mixed and Multiple Research Designs
  • Planned or Emergent Design: A mixed or multiple research design may be planned up front as guided by the research problem or question, or it may organically evolve over time because a first approach may not work as well as anticipated (for example, due to low participation), or it may lead to insights that beg for confirmation, expansion, explanation, or triangulation.
  • Inductive/Deductive Dominance: An MMMR design typically involves both inductive (from data to theory) and deductive research (from theory to data), but one approach usually dominates or takes priority over the other driven by the philosophical world view of the researcher or by the nature of the research problem or question. Researchers often conflate inductive and deductive research with quaL and quaN methods, respectively, but Young et al. explain how inductive and deductive research may fall along a “spectrum” of quaL and quaN data young_spectrum_2020 .
  • Timing of Procedures: The various procedures may be conducted in sequence or in parallel. A sequential design is driven by the emergence of new or refined research questions, while a parallel approach may be motivated by the expertise of the research team, access to research participants and data, or timing constraints.
  • Mixed or Multi Method Design: When the research mixes quaL and quaN data (during collection or analysis), the design is referred to as mixed rather than a multi methods design, which uses both quaL or quaN procedures.
  • Data Integration: How the data is integrated during collection or analysis shapes the research design and the research insights ladner_mixed_2019 (as mentioned above). For example, in the case of parallel MMMR, quaN and quaL data may be merged to lead to the findings or insights, or in the case of a sequential study, one type of data may build on top of or add to another type (see Figure 1).
4.2 Research Designs
The characteristics mentioned above and how they are composed lead to a constellation of different research designs. We use the nomenclature for research designs as proposed by Creswell and Plano Clark creswell_best_2011 as we describe some popular design variations in the following.
  • Exploratory Sequential: This sequential design often prioritizes a quaL method (normally the dominant method) followed by a quaN procedure that builds on insights from the first procedure. The design may be planned up front or it may emerge as insights are uncovered from the first procedure.
  • Explanatory Sequential: This sequential design starts with and prioritizes a deductive quaN method (normally the dominant method) that is followed by a quaL method to help explain the insights from the first procedure. The design may be planned up front or the need for the second procedure may emerge, especially in light of some unexpected insights or where some insights need further interpretation or explanation.
  • Convergent Parallel: The convergent parallel design involves both a quaL procedure and a quaN procedure to bring complementary insights on a research question. The goal may be to uncover both convergent and divergent insights from both procedures that are planned in parallel.
  • Embedded: An embedded or nested design is when one procedure is dominant (quaL or quaN ) but another procedure appears as a “strand” within the first procedure. For example, as part of a quaN survey, a researcher may choose to interview select participants while carrying out the survey to gain more quaL insights on their answers.
  • Multi-method: A multi-method design may be planned or emergent and involves either two or more quaN procedures or two or more quaL procedures, but not a mix of both quaL and quaN. The main goal is typically to expand on the insights across procedures, and they may be carried out sequentially or in parallel.
5 Researcher Personas and MMMR Design Scenarios
To illustrate the principles, characteristics and MMMR designs we presented so far in this paper, we showcase four fictitious personas as guiding examples. These four fictional software engineering researchers are at different stages in their research careers, pursuing quite different research projects, each with a different reason why a mixed or multi methods design is appropriate. We describe these below and mention the decisions and challenges they faced in conducting their research. For the four scenarios, we provide illustrations of the research designs (see Figures 3, 4, 5 and 6). Later we summarize the four scenarios in Table 1.
5.1 Ali’s Study: The Impact of Security Experts Assigned to Development Teams
Figure 3: Ali follows an exploratory sequential design, studying the impact of security experts on development practices. The unit of analysis is practices as informed by developers across both procedures.
Ali (he, him), is a postdoc studying the introduction of a new process whereby security experts are assigned to development teams to improve the compliance of security practices by developers at a large company. Ali’s research question is to ask if this change impacts developer motivations and capabilities for writing more secure code.
As this is a new process and there is little research to build on, Ali first conducts an inductive exploratory open-ended interview with a small number of developers at a company to find out how this change has impacted developer motivations and behaviours for writing more secure code. As Ali conducted these interviews, he took care to protect the identity of the team members he interviewed, with both the security and ethical concerns in mind. The findings from these interviews help identify a number of concepts that help describe changes in motivation.
Ali wishes to know if the interview findings from this team hold (i.e., generalize) across a broader set of developers. He expands the preliminary exploratory approach and uses the insights from the initial interviews to develop a follow-up closed question validating survey with a larger population of developers beyond the company to confirm if positive changes in motivation occur when security experts are added to software teams. Again, the identity of the developers was protected and care was taken to ensure the presentation of the findings neither compromises the developers’ roles nor the confidentiality of the projects they work on.
Both methods were conducted in a rigorous way and led to novel insights about the impact of security experts on developer teams and how those hold across a larger population of developers. Ali planned to have both studies at the start of the research and sees the qualitative insights as the dominant contribution in this exploratory sequential research design. The tradeoff Ali made in this design is to focus some of his research effort on the follow-up survey to generalize the findings over achieving potentially deeper insights from longer or more interviews in the first phase.
5.2 Sam’s Study: The Impact of a New Code Review Tool on Code Quality
Figure 4: Sam follows an explanatory sequential design to study a new code review tool. The unit of analysis is telemetry process metrics in (1), while in (2) it is perception of the process informed by developers.
Sam (they, them) is an early stage Ph.D. student collaborating with a small company that is trialing a new tool to automate part of the code review process with hypothesized improvements on time-to-merge of code changes and on the quality of the reviewed code. They start their research with a deductive quantitative data mining study to investigate their hypothesis that this new tool (before and after use) improves time-to-merge and reduces the number of bugs captured in their industry partner’s code repository. As they mine the data, Sam ensures that their findings do not reveal whose coding or code review activities may have resulted in the future introduction of bugs with other team members or leads.
From the mining study phase, Sam finds that the tool results in the hypothesized faster time-to-merge but not a lower number of bugs after merge. This insight was not expected, thus to understand why, they then conduct open ended interviews with a select number of developers to explore how the tool was used and to investigate inductively why it did not have the hypothesized impact on bug count. Again, care is taken to protect the identity of the developers interviewed, but also to ensure the interviews did not make them feel stressed given the presence of some bugs they may have introduced. From the interviews as well as the mined data, novel insights about the tool emerged.
The use of mixed methods was not intended up front, but was an emergent or organic research design decision initiated by the surprising findings. These integrated insights can be used to shape code review tool design accordingly. The resulting design is referred to as an explanatory sequential design as mentioned above (see Figure  4). A tradeoff was made to halt further mining of the telemetry data and instead to conduct interviews with developers to explain the surprising results, but also to shed light on more actionable insights for how the code review tool could be improved.
5.3 Vicki’s Study: The Adoption of Generative AI at a Software Company
Figure 5: Vicki follows an emergent convergent parallel design to study the adoption of Generative AI. The unit of analysis is code changes by directly observing code.
Vicki (she, her) is an experienced industry researcher working at a software company, and her research aim is to understand the grassroots adoption of generative AI technologies across her company. As this technology is so new, and not much is known about how it can and will be used to enhance developer productivity, she decides to inductively consider the AI adoption from two angles in parallel. First, Vicki uses an ethnographic approach of observations to glean how the generative AI tool is used by developers at her organization. She observes its adoption on her own team of developers, while taking care of ethical concerns by anonymizing insights on individual productivity as the use of a new tool like AI may stress developers if they worry about losing their jobs.
In parallel to complement these observations, she carefully mines logging data from the AI tool used by multiple teams and in projects at her organization to understand more about the scale of adoption and how the tool is influencing developer productivity in terms of number of commits and speed of feature development. She recognizes that mining this data comes with the potential risk of introducing privacy concerns as she works at the company, but she mines in a way to obfuscate developers’ identities and does this to confirm her observations (which could be biased by her own impressions).
The use of a convergent parallel design that involves conducting both procedures in parallel provides complementary and novel insights on the adoption of this new potentially disruptive technology at her company.
5.4 Zara’s Study: Understanding the Effectiveness of an Automated tool for Generating Tests
Figure 6: Zara follows multi-method design to study a test generation tool. The unit of analysis is the developers in (1), while in (2) it is the test cases generated by the tool.
Zara (she, her) is a late stage Ph.D. student who aims to investigate how effective an automated tool she designed is for generating tests. Before asking practitioners, she initially conducts a lab experiment with senior student developers, where half use the tool and half do not. She takes care during recruitment (via a university professor) to be cautious about how she recruits the students.
She starts deductively with the hypothesis that the tool will help with faster test generation. She assigns the same testing tasks to both groups. Through self reports from the developers during the experiment, Zara finds that the student developers report that through the use of her tool they are faster at generating test cases, supporting her hypothesis and also revealing that they believe the tool makes them faster.
This positive finding encourages her to expand her research question to investigate if the tests generated by her tool are as or more effective at finding bugs. She follows up by mining test and bug data generated during the lab experiment to investigate if the tests generated are effective at defect localization. Although she is aware of the identity of the students from the experiment, she takes care to obfuscate their data when she mines it.
Note she expands not only the research question she is asking, but she also changes the unit of observation in her study (from observing developers to analyzing the test cases generated). In this case, her research design is a multiple methods design rather than a mixed design as she does not mix qualitative and quantitative data while both of her studies are deductive by nature. Following her study, she has gained novel insights on how students may benefit from the test generation tool. An alternative design to this multiple methods design could have been to only mine data from the students’ activities. However, she used two methods as she was also interested to gain insights on their perception of the tool’s effectiveness.
Type of Design Design flow Rationale(s) Persona
Exploratory Sequential
QuaL: Inductive Open-ended Interviews $\rightarrow$
QuaN: Deductive Closed Survey
Expansion
Triangulation
Development
1. Ali
Security
Compliance
Explanatory Sequential
QuaN: Deductive Data Mining $\rightarrow$
QuaL: Inductive Interviews
Initiating
2. Sam
Code review
Convergent Parallel
QuaL: Inductive Ethnographic Observations $||$
QuaN: Inductive Data Mining
Complementary
3. Vicki
Adoption of
Generative AI
Multi-Method Design
QuaN: Deductive Lab Experiment $\rightarrow$
QuaN: Deductive Data mining
Expansion
4. Zara
Test generation
Table 1: The first three scenarios are mixed method designs, while the fourth is a multiple methods design. Note that there may be several reasons for conducting a mixed or multiple method design, and this may not be apparent up front (e.g., in the case of Sam).
5.5 Learning from Other Examples
So far, we have shared that using MMMR enables SE researchers to study their topics from a more holistic perspective, expanding and complimenting methods to achieve more and novel insights. Furthermore, we provide personas and scenarios to demonstrate the broad and varied landscape of possible MMMR designs in SE research. The ACM SigSoft Empirical Standards for Multiple and Mixed Methods also provides a list of example MMMR papers sigsoft , but care has to be taken in using these and other examples of MMMR designs as “exemplars” for future work. With this in mind, we next present “antipatterns” to watch out for when designing or critiquing an MMMR design. An earlier version of these antipatterns also appears in the same Standards, contributed by the first author of this paper.
6 Antipatterns of Mixed and Multi Method Research Designs
The principles and research design patterns we described above can be considered as useful guidelines to follow when designing a study. And through the scenarios above, we illustrated how the principles of MMMR could be followed. In the following, we provide examples of “what not to do” and “what to avoid”.
As researchers and reviewers, we have noted a number of common antipatterns in how mixed methods are used or presented. Some of the antipatterns we encountered originate in issues with how the research is presented (rather than inherent to the conduct of the research). We introduce the presentation antipatterns first noting how they can be potentially fixed. The other study design antipatterns are harder to “fix” after the research has been conducted, but hopefully can be spotted at design time (or during review of a paper). We recommend researchers think about these antipatterns early in their research process.
6.1 Presentation Antipatterns
  • Uninvited guest or party crasher: When one of the methods used is not clearly introduced in a paper introduction or methodology and makes an unexpected entrance in the later part of the paper, such as in the discussion or limitations sections. To fix this, the author should clearly mention all methods in the introduction and methodology sections of the paper, justifying why they were used. Even with emergent designs, as in the cases of Sam and Zara above, all methods used should be presented early in the paper with a mention that the design was emergent.
  • Smoke and mirrors: This antipattern can be spotted when a paper attempts to oversell a study as an MMMR when one of the approaches merely offers a token or anecdotal contribution to the research motivation or findings. This can be fixed by clarifying the presentation in the paper.
  • Limitation shirker: This antipattern occurs when there is a failure to discuss limitations from all methods used or from their integration in a study report. MMMR will have a much longer and more in depth limitation section than a single study paper, and the limitations reported should be congruent with the procedures used or mixed.
6.2 Study Design Antipatterns
  • Missing the mark: Misalignment of MMMR design with the research question/objective. This can be seen when one (or even both) of the methods do not make sense in light of the research question being asked. The fix for this should happen at the study design phase by asking if every method answers the core research questions being asked (whether planned up front or emergent). Note that sometimes the unit of observation may be changed but the unit of analysis remains the same. For example, Zara switched to looking at telemetry data rather than observing developers, but her unit of analysis remained the developers’ use of the tool for test generation.
  • Selling your soul: This antipattern is related to “Missing the Mark” and may occur when a researcher employs an additional method to appeal to an anticipated reviewer who has preference for a particular methodology, but the use of the procedure does not contribute substantively to the research findings. This can be addressed by using the method one believes in or by ensuring that the researcher draws out meta-inferences from the multiple procedures used.
  • Cargo cult research [^6] Cargo Cult can be a metaphor for superficial imitation of a process without basic understanding of its mechanism.: Using methods in which the research team lacks expertise but that they have observed being used to generate “exciting findings” in well cited papers. Collaborating with other researchers with broader expertise may help avoid this antipattern.
  • Sample contamination: This antipattern may occur in an MMMR sequential design where the same participants are used in multiple, sequential methods without accounting for potential contamination from earlier method(s) to later ones (where participants have already seen an intervention or know about the research). However, in some cases the design itself may involve selecting participants (e.g., for interviews) from an earlier method (e.g., a survey) and the resulting “contamination” is intentional. Any possible contamination should be justified or reported, and potentially mentioned in limitations.
  • Lost Opportunity: In an MMMR sequential design, when there is a failure to use findings from an earlier study during the development of an instrument for a study in a later phase of the research. This is an issue to be caught during the implementation of the study as it cannot be fixed through the presentation (alternatively, the studies can be written up as a two-phase study or two separate studies with two different contributions).
  • Integration failure: We talked about failure to integrate data or insights as a quality attribute earlier in the paper. This antipattern occurs when there is poor integration of findings from all methods used in the study design. This can be fixed by planning how the results/findings will be integrated and making sure that the results/findings from both studies are integrated or merged in the results/findings section of the paper.
  • Questionable Ethics: Finally, even when all of the above antipatterns are avoided, i.e., the MMMR study is applied and presented well, it may come across as having questionable aims, or as using questionable means to achieve the aims. The researcher should clearly state the motivation for the study and expected benefit to society or research to help justify the possible risks a study may involve.
7 Conclusions
In this paper, we provide an overview of how mixed and multi methods research designs can be used in software engineering, with the intention to inspire their use. Other resources, such as the ACM Sigsoft Empirical Standards sigsoft , as well as our earlier paper on selecting a research method in software engineering easterbrook_selecting_2008 , only lightly touch on the concepts and challenges of using mixed methods research in software engineering. As can be seen from Section 2, there is a long history of research behind the mixed method paradigm. Deciding what to include (and exclude) in this paper from this large body of research was challenging and we focused on the principles, characteristics and designs that we hope will support software engineering researchers in applying mixed methods to lead to impactful insights. Still, applying mixed methods is not a trivial exercise, as we try to demonstrate through the identification of many possible antipatterns.
The “selecting empirical research methods” paper easterbrook_selecting_2008 notably elaborates on topics such as theoretical stance and the role of theory in quantitative, qualitative and mixed research. The main takeaways from that paper also apply here: a mixed or multi methods paper should have a clear research question that aligns with the research design; the procedures used should answer this question; the limitations of any one method or procedure can be used to motivate the choice of other methods, but the limitations of and threats associated with mixing should also be reported; and the role of theory and whether the research is inductive or deductive should drive the research design and shape the integrated findings reported.
We also elaborate on four core principles of mixed and multi methods research that are: providing methodological rationale for the study design; ensuring the design leads to novel insights by the mixing of methods; assuring the rigorous use of individual methods and that they are mixed in a rigorous way; and finally, conducting the study in an ethical manner. These principles, at a high level, also apply to any research study, but we elaborate on how these principles can be considered specifically within the context of a mixed or multi methods research study.
We further describe the core characteristics that shape a diverse landscape of mixed or multi methods research designs. These characteristics, which include the timing of procedures, the inductive or deductive nature of the research, the research design planning or emergence, how data is integrated, and the mixed or multi design, should be carefully considered up front and throughout the research process, and presented clearly in any research manuscripts. We do not provide a comprehensive description of all possible designs, but through several personas and research scenarios, we show how the principles and characteristics can be considered as tradeoffs and design choices are made. The integration process throughout the research design is what sets mixed methods apart.
Our paper further suggests a number of practical considerations. For example, how to visualize the mixed or multi methods research design (as we show for the four personas), some potential pitfalls to try to avoid (articulated as research antipatterns), and which limitations to consider and present (guided in part by the principles). Mixed or multi methods, as mentioned, can be used to offset the limitations of a single method, and can be powerful for investigating socio-technical phenomena in software engineering, but they are challenging to use and often require broad expertise and substantial time commitments from the researchers involved.
This paper is a first step in understanding how to use mixed methods in software engineering research and we hope that the SE research community will find this guide helpful. In closing, we suggest researchers adopt an open and creative stance to research design, and to share back with the community their insights gained not only on software engineering, but also on their research designs as they use mixed or multi methods.
Appendix: Glossary
Research Paradigms
A research philosophy refers to beliefs, assumptions, values, and practices constituting a particular worldview or framework for understanding phenomena within a specific discipline or field of study (e.g., Positivism, Constructivism, Pragmatism, etc.). Paradigms provide a lens through which researchers interpret and make sense of reality, guiding their observations, theories, methods, and interpretations.
Strategy
A research strategy refers to the overarching plan or approach that guides the conduct of a research study (e.g., ethnography, socio-technical grounded theory, case study, action research, survey, experiment, etc.). It outlines the steps, methods, and procedures that will be employed to address the research question or objectives effectively.
Research Design
A research design involves selecting the most appropriate research design based on the nature of the research question and the desired outcomes (i.e., the researcher’s methodological choice if using the mono-method, multi-method, or mixed methods).
Data Collection Method
Researchers must decide how to gather data relevant to their research question (i.e., techniques and procedures). This may involve methods such as surveys, interviews, observations, or a combination of these approaches.
Data Analysis Techniques
Researchers specify the methods for analyzing the data collected during the study. These may involve statistical techniques, socio-technical grounded theory (STGT) for data analysis hoda_socio-technical_2021 , thematic analysis braun2012thematic , or other analytical approaches.
Qualitative and Quantitative
Qualitative (quaL) and quantitative (quaN) are two primary approaches to research, each with distinct characteristics. In this paper, we refer to quaL data as text and image-based evidence (e.g., words used to describe experiences or stories). In contrast, quaN data relates to numeric-based evidence (e.g., device analytics or Net Promoter Score).
Inductive and Deductive
Inductive and deductive reasoning are two complementary approaches used in research to develop theories, hypotheses, or generalizations. Inductive reasoning involves moving from specific observations or patterns to broader generalizations
Deductive reasoning involves moving from general principles or theories to predictions or hypotheses about specific instances. While inductive is commonly associated with quaL research methods, deductive is frequently associated with quaN research methods. However, this can vary, as exemplified in our scenarios (Section 5).
Meta-inferences
Meta-inferences refer to making inferences or drawing conclusions about integrating the quaLand quaNmethods. When integrating quaLand quaNdata is a central element of the MMMR, how the researcher represents this integration and the interpretation of these findings becomes critical. During the MMMR discussion, the meta-inferences depicted during this process should be articulated to present how the use of an MMMR approach advanced a great understanding of the substantive topic compared to a mono-method approach.